OpenStudy (anonymous):

how do you graph x= y^2 - 9 on the calculator?

OpenStudy (eyust707):

what kind of calculator do you have?

OpenStudy (amistre64):

by typing in y = x^2-9 and leaning it over on its side

OpenStudy (eyust707):

hahaha i was gunna say the same thing

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

just any graphing calculator

OpenStudy (amistre64):

or split it on 2 ys y1 = sqrt(x^2+9) y2 = -sqrt(x^2+9)

OpenStudy (anonymous):

so what's the x and y intercepts ?

OpenStudy (anonymous):

-3, 3

OpenStudy (amistre64):

its a parabola that on its side ....

OpenStudy (amistre64):



OpenStudy (anonymous):

thanks